Parth (parthkohli):

Mechanics 1+2 question.

Parth (parthkohli):

A load of 981 N is suspended from a steel wire of radius 1 mm. What is the maximum angle through which the wire with the load can be deflected so that it does not break when the load passes through the equilibrium position? Breaking stress is \(7.85 \times 10^8 N/m^2\).

Parth (parthkohli):

This will test your knowledge of mechanics 1 (rigid body dynamics) and mechanics 2 (fluids and bulk matter)

ganeshie8 (ganeshie8):

stress = Y * strain

ganeshie8 (ganeshie8):

\[\dfrac{F}{A} = Y*\dfrac{\Delta L}{L}\]

ganeshie8 (ganeshie8):

just need to express \(\Delta L\) in terms of angle

Parth (parthkohli):

Hmm, that's not it.

ganeshie8 (ganeshie8):

i don't see how this can be anything different... let me attempt :)

OpenStudy (anonymous):

the answer is 3 XD. You can medal me your welcome

ganeshie8 (ganeshie8):

I think 981 N + centrifugal force cannot exceed the maximum stress

ganeshie8 (ganeshie8):

If that is the case, do i get : \[981 +ma_c = 7.85*10^8*\pi*10^{-6} \] ?

ganeshie8 (ganeshie8):

somehow i need to express \(a_c\) as a function of the deflection angle

Parth (parthkohli):

You're getting close.

ganeshie8 (ganeshie8):

let me grab pen and paper

ganeshie8 (ganeshie8):

after some rearrangement im getting \[981(1+ h) = 7.85*10^8*\pi*10^{-6} \]

ganeshie8 (ganeshie8):

need to express \(h\) in terms of the deflection angle

ganeshie8 (ganeshie8):

but \(h\) depends on the length of wire too so above method should lead me to a dead end hmm

OpenStudy (baru):

at equilibrium, potential energy energy = 0 we have 0.5mv^2=mg h cos( ) we could eliminate h using that

Parth (parthkohli):

Yeah, exactly. You need to employ a concept in rigid-body mechanics. Baru is getting there!

OpenStudy (baru):

981 + \(mv^2/r\) = stress * area

ganeshie8 (ganeshie8):

still wondering.. not getting a clue with that, i think baru and me are in same boat :)

OpenStudy (baru):

r=h

ganeshie8 (ganeshie8):

brb

OpenStudy (baru):

981 + mv^2/h = stress * area 0.5mv^2=mgh cos( ) substitute for v^2 in eq 1, h will get cancelled solve for cos ( ) ?

Parth (parthkohli):

Yeah, hint: vertical circular motion.

Parth (parthkohli):

Should I spoil it now?

ganeshie8 (ganeshie8):

nope, i got it

Parth (parthkohli):

Oh, how?

ganeshie8 (ganeshie8):

eating now, will post in 10 mnts

Parth (parthkohli):

:)

ganeshie8 (ganeshie8):

Nice ! thanks to baru for hinting conservation of energy

OpenStudy (baru):

Sure :) Still don't get it though :(

Parth (parthkohli):

\[\frac{mv^2}{2}=mg\ell (1 -\cos \theta_{max})\]\[\Rightarrow \frac{mv^2}{\ell} = 2mg(1-\cos \theta_{max})\]Now\[T_{breaking}-mg = \frac{mv^2}{\ell}\]\[\Rightarrow T_{breaking} = 3mg - 2mg \cos \theta_{max} \tag{using previous equation}\]Now use\[T_{breaking}=\rm{area \times breaking ~stress}\]

OpenStudy (baru):

Oh ok..Potential energy is mgh(1-cos)....I missed that

ganeshie8 (ganeshie8):